OpenStudy (kaiiskool):

John has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $26.70. Let x represent the number of quarters in the jar. Which equation can be used to find the number of quarters?

OpenStudy (anonymous):

do you have any choices to choose from?

OpenStudy (kaiiskool):

0.05x + 0.25(150 – x) = 26.70 0.25x + 0.05(150 – x) = 26.70 0.25x + 0.05(x – 150) = 26.70 0.05x + 0.25(x – 150) = 26.70

OpenStudy (anonymous):

0.25x + 0.05(150 – x) = 26.70

OpenStudy (kaiiskool):

Thank You so much! my brain is off today.

OpenStudy (anonymous):

your welcome ^.^